OpenStudy (anonymous):

Light travels in a vacuum at a speed of roughly 2.998x10^8 m/s and its speed in air is only negligible (0.03%) slower than that. a)How long does it take light to traverse 1.0 ft? b) That means when you look at something 1000 m away, you are seeing it how many seconds back in times?

OpenStudy (anonymous):

a. C=3X 10^8 m/s; d= 1 ft X 1m/3.3 ft = 3X 10^-1 m ; V= d/t solve for t ; t= d/V V=c for light. t= 3X 10^-1 m/ C=3X 10^8 m/s= 1X 10^-9 s for part b.. do the same thing.